OpenStudy (anonymous):

the gasoline gauge on a van initially read 1/4 full. When 10 gallons were added to the tank the gauge read2/3 full. How many more gallons are needed to fill the tank?

OpenStudy (anonymous):

1/4 + 2/3 = 3/12 + 8/12 = 11/12 Leaving 1/12 left to fill the tank. I think....not 100% sure

OpenStudy (anonymous):

am I way off mathstudent ?

OpenStudy (mathstudent55):

This is a little tricky.

OpenStudy (anonymous):

1/4 + 2/3 + 1/12 = 3/12 + 8/12 + 1/12 = 12/12 (full tank)

OpenStudy (mathstudent55):

Adding 10 gallons of gas is not adding 2/3 of the tank to teh 1/4 tank that was already there.

OpenStudy (anonymous):

yes this is very confusing to me!

OpenStudy (mathstudent55):

Adding 10 gallons made the tank 2/3 full.

OpenStudy (anonymous):

10 gallons fills 5/12 up. Does that make sense. 2/3 - 1/4 = 8/12 - 3/12 = 5/12

OpenStudy (mathstudent55):

OK, think of it like this. The full tank takes x gallons.

OpenStudy (mathstudent55):

The tank was reading 1/4 full. That means the tank has (1/4)x gallons, or 0.25x gallons.

OpenStudy (mathstudent55):

Then 10 gallons were added. That means 0.25x + 10

OpenStudy (mathstudent55):

After adding the 10 gallons, the tank reads 2/3 full. This is (2/3)x, so 0.25x + 10 = (2/3)x

OpenStudy (anonymous):

10 gallons is 5/12, so 20 gallons is 10/12 ...right ?

OpenStudy (mathstudent55):

yes

OpenStudy (mathstudent55):

(1/4)x + 10 = (2/3)x Multiply both sides by 12 to get rid of parentheses: 3x + 120 = 8x

OpenStudy (mathstudent55):

subtract 3x from both sides 120 = 5x

OpenStudy (mathstudent55):

Divide both sides by 5: x = 24

OpenStudy (mathstudent55):

The fill tanks has a 24-gallon capacity.

OpenStudy (anonymous):

you are so smart.....I am jealous :)

OpenStudy (mathstudent55):

Since the tanks became 2/3 full by adding the 10 gallons, it still needs 1/3 of its capacity to be full. 1/3 of 24 gallons = 8 gallons

OpenStudy (mathstudent55):

Now le's check 24 gallons to see if it's correct.

OpenStudy (anonymous):

1/4(24) + 10 = 2/3(24) 8 + 10 = 16 ....are you sure

OpenStudy (mathstudent55):

The tank was 1/4 full. 1/4 of 24 gal is 6 gal. Then 10 gal was added. Now the tanks has 6 gal + 10 gal = 16 gal This is 2/3 of the tank. 16/24 = 2/3, so it's correct. The answer is 8 gallons.

OpenStudy (mathstudent55):

1/4 0f 24 = 6 1/3 of 24 = 8

OpenStudy (anonymous):

oops...I did it wrong...it does equal

OpenStudy (mathstudent55):

We use both 1/4 tank in the beginning of the problem, and 1/3 tank at the end.

OpenStudy (anonymous):

you should be a teacher

OpenStudy (mathstudent55):

thanks